ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 35L50, 35Q35, 73E50 Hayes, Brian and Schaeffer, David, Duke University, Durham, NC, USA Stress-Controlled and Velocity-Controlled Shear Waves in a Saturated Granular Medium We study the interaction of plane shear waves, when either the stress or velocity is a given periodic function on the boundary. The soil is modelled as a continuum, obeying a hypoplastic flow rule. In both cases, wave interactions are confined to a strip of width x_{I}, along the boundary. As t → ∞, x > x_{I}, the velocity and stress approach asymptotic, time-independent values. For the stress-controlled case, the saturated state depends smoothly on the boundary data. In the velocity-controlled case, however, the asymptotic state exhibits a wildly discontinuous, fractal behavior, as the boundary data is varied.